A car radiator holds 2 gallons of fluid and initially is empty. If a mixture of water and antifreeze contains 70% antifreeze, how much of each should be combined to fill the radiator with a 50% antifreeze mixture?
Answer by rwm(914)

from the antifreeze angle
.7x=.5*2
x=1.43
so add 2-1.43 = .57 gallons of water.
from the water angle
70% antifreeze is 30% water
.30x+y=.5*2
x+y=2
x=1.42857 gallons antifreeze , y=0.571429 gallons water